Solve the cubic equations:
a) (x - 3)^2 = 4/x
b) x(x+3)(x-1) = x + 8
c) x^3 + 4 = x(x + 4)

Answer by ewatrrr(24785)


Hi,
a) (x - 3)^2 = 4/x OR x^3 -6x^2 + 9x - 4 = 0 where x = 1 is one solution
Applying division process:
____________________
x -1 |x^3 -6x^2 + 9x - 4 = x^2 - 5x + 4
(x-1)(x-4)(x-1) = 0 x = {1,4}
b) x(x+3)(x-1) = x + 8 OR x^3 + 2x^2 -4x -8= 0 where x= 2 is one solution
Applying division process:
___________________
x -2 |x^3 + 2x^2 -4x -8 = x^2 - 4
(x-2)(x-2)(x+2) = 0 x = { 2, -2}
c) x^3 + 4 = x(x + 4) OR x^3 - x^2 - 4x + 4 where x = 1 is one solution
Applying division process:
___________________
x -1 |x^3 - x^2 - 4x + 4 = x^2 - 4
(x-1)(x-2)(x+2) x = {1,2,-2}